Intel® Core™ i5-7500 is one of Intel's 7th generation of microprocessors and was launched earlier in 2017 and has been used in many computers, including Dell OptiPlex 7050 Tower.

a) How many CPU does this microprocessor have? What's its cache memory capacity?

b) According to Intel, this microprocessor's max clock speed can reach 3.8 GHz. If an instruction requires 5 clock cycles to complete execution in this microprocessor, then, in terms of nanoseconds, how long does it take to get executed? Round the number of your answer to the nearest thousandth (e.g., 28.157 or 0.540).

2. Assuming you have an 10-year-old desktop PC that is connected to the Internet via a cable using the Ethernet port. It has no built-in Wi-Fi capability, that is, it has no wireless Internet connection. Based on your experience or learning of this chapter or additional research, provide two approaches to get wireless connectivity on your old PC.
Answer #1

a) Intel® Core™ i5-7500 has 4 physical cores and 4 threads. The base clock frequency of each core is 3.4 GHz and the turbo frequency is 3.8 GHz.

The processor has 6 MB smart cache.

b) Processor speed is 3.8 GHz

So, in one second 3.8 * 109 instruction cycles can be executed.

The processor has four cores which means that at a time, four instructions can be fetched and executed simultaneously.

Each core has 3.8 GHz speed and all the four cores can execute four instructions at a single instruction cycle execution time.

For 4 instruction cycle the time required is (1/(3.8*109)) seconds = 0.263 nanoseconds

For additional 1 clock cycle the time required is (1/(3.8*109)) seconds = 0.263 nanoseconds

Total time required is (0.263 + 0.263) nanoseconds = 0.526 nanoseconds

2) Two approaches to get wireless connectivity on a ten years old computer are :

(i) Using a wireless dongle : There are a lot of high speed wireless dongles available in the market. This can provide mobility and wireless connectivity.

(ii) Using a wireless router : Router can provide wireless connection within a specific range. The old computer can connect to the router which can provide wireless connection.

Hope this helps.